alexSalvadori83:

A store manager timed Janette to see how long it would take her to fold and put away a sweater, a shirt, a pair of pants, and a scarf. It took her 26.1 seconds for the shirt, 24.3 seconds for the sweater, 32.8 seconds for the pants, and 18.2 seconds for the scarf. What was the average time it took Janette to fold and put away all four items?

Extrinix:

So to find the average, you add all of the timings together and divide by the amount of timings. \(\dfrac{26.1+24.3+32.8+18.2}{4}=?\) \(\dfrac{101.4}{4}=?\) So you can solve from there.